# Test-Time Anchoring for Discrete Diffusion Posterior Sampling

### v1

@@ B.1.2 Reference-based Stylization
For each reverse step, we perform $100$ optimization steps using an Adam optimizer with a learning rate of $0.1$. Gradients from this style loss are backpropagated through the frozen VQ-VAE decoder ${\mathcal{D}}(\cdot)$ and the straight-through estimator $\tilde{{\mathbf{x}}}$ to update all the entries of the conditional probability table by directly updating the logits $\varphi_{t(i)}=\{\varphi_{t(i)}^{l}\}_{l=1}^{L}$.
In the second stage (§3.2.2), the posterior estimate ${\mathbf{x}}_{\varphi}({\mathbf{z}}_{t(i)})$ obtained from the first stage is used adaptively unmask anchor tokens in the sequence. Both the processes continue over $15$ total steps following a cosine mask schedule. The fully unmasked sequence satisfies the stylistic constraints without requiring any task-specific retraining of the foundation model.
B.1.3 High-Resolution Inference
@@
B.2 Algorithm Details
1:Input: Diffusion steps $T$, measurement $\mathbf{y}$}, denoiser ${\mathbf{x}}_{\theta}^{\mathrm{logits}}(\cdot)$, operator ${\mathcal{A}}(\cdot)$, decoder ${\mathcal{D}}(\cdot)$, Codebook ${\mathcal{C}}=\{{\mathbf{c}}_{k}\}_{k=1}^{K}$, Lookup-Free Quantizer ${\mathcal{Q}}_{\mathrm{lfq}}$.
@@ B.2 Algorithm Details
5:for $i=1$ to $T$ do
6:  $\varphi\leftarrow{\mathbf{x}}_{\theta}^{\mathrm{logits}}({\mathbf{z}}_{t(i)})$ $\triangleright$ Quantized Expectation §3.2.1
7:  ${\mathbf{x}}_{\theta}({\mathbf{z}}_{t(i)})=\mathrm{Softmax}({\mathbf{x}}_{\theta}^{\mathrm{logits}}({\mathbf{z}}_{t(i)})$
8:  for $m=1$ to $M$ do
9:   ${\mathbf{x}}_{\varphi}({\mathbf{z}}_{t(i)})=\mathrm{Softmax}(\varphi)$
10:   $\bar{{\mathbf{x}}}\leftarrow\sum_{k=1}^{K}{\mathbf{c}}_{k}{\mathbf{x}}_{\varphi}({\mathbf{z}}_{t(i)})$
11:   ${\mathbf{x}}={\mathcal{Q}}_{\mathrm{lfq}}(\bar{{\mathbf{x}}})$
12:   $\tilde{{\mathbf{x}}}\leftarrow\bar{{\mathbf{x}}}+\big[{\mathbf{x}}-\bar{{\mathbf{x}}}\big]_{\mathrm{sg}}$
@@
14:   $\mathcal{L}\leftarrow\mathcal{L}_{\mathrm{measurement}}({\mathcal{A}}(\hat{{\mathbf{x}}}),{\mathbf{y}})+\lambda_{\mathrm{p}}\mathcal{L}_{\mathrm{perceptual}}({\mathcal{A}}(\hat{{\mathbf{x}}}),{\mathbf{y}})+\lambda_{\mathrm{pp}}{\mathcal{L}}_{\mathrm{prior}}({\mathbf{x}}_{\varphi}({\mathbf{z}}_{t(i)}),{\mathbf{x}}_{\theta}({\mathbf{z}}_{t(i)}))$
@@
16:  end for
17:  ${\mathbf{x}}_{\varphi}^{*}({\mathbf{z}}_{t(i)})=\mathrm{Softmax}(\varphi)$ $\triangleright$ Anchored Remasking §3.2.2
18:  $\bar{{\mathbf{x}}}^{*}\leftarrow\sum_{k=1}^{K}{\mathbf{c}}_{k}{\mathbf{x}}_{\varphi}^{*}({\mathbf{z}}_{t(i)})$
19:  ${\mathbf{x}}={\mathcal{Q}}_{\mathrm{lfq}}(\bar{{\mathbf{x}}}^{*})$
20:  $\kappa^{l}\leftarrow\langle({\mathbf{x}}_{\varphi}^{*}({\mathbf{z}}_{t(i)})^{l},{\mathbf{x}}^{l}\rangle$ for $l=1,\dots,L$
21:  ${\mathcal{P}}_{t(i)}\leftarrow\{l:\kappa^{l}\geq\tau_{t(i)}\}$
22:  ${\mathbf{z}}_{s(i)}\leftarrow\mathrm{UpdateState}({\mathbf{z}}_{t(i)},{\mathbf{x}},{\mathcal{P}}_{t(i)})$
23:end for
@@
25:return $\hat{{\mathbf{x}}}_{\mathrm{final}}$
Algorithm 1 Test-Time Anchored Posterior Sampling (APS)
Algorithm 1 details our APS procedure. The process begins with a fully masked latent space, ${\mathbf{z}}_{1}$, and iteratively refines the image over $T$ reverse diffusion steps. Each step features an inner optimization phase designed to align the model’s predictions with the given measurement $\mathbf{y}$}. To enable gradient-based optimization through the discrete quantization step which assigns each dimension of an embedding to its nearest value in $\{-1,1\}$ we employ the straight-through estimator (STE) via the stop-gradient operator $\big[\cdot\big]_{\mathrm{sg}}$.
This optimization is guided by a composite loss function. For reconstruction, we primarily use the Mean Absolute Error (MAE, or L1 loss), which we find produces perceptually superior results compared to the Mean Squared Error (MSE). While MSE corresponds to maximizing the Gaussian log-likelihood, MAE is typically more robust. This is supplemented with a VGG perceptual loss to enforce similarity in the features extracted from the measurements, further improving visual quality in the image space [^4] The LPIPS metric measures visual quality in the image space.. Once the optimization at a given step is complete, the Anchored Remasking strategy uses the tilted measure to estimate confidence assigned to the token chosen via quantization. This is followed for each position to selectively unmask tokens for the next iteration.
@@ B.3 Ablation Study
Figure 6: Ablation study of design choices in APS. We study the posterior sampling problem with the input of the noisy image of a bird (top row, third image from left). In the base case of using MMada (second row), since it does not take an image as an input (but only text), the final image is inconsistent with the noisy input image (image of person instead of bird). Thus, this shows the effect of standard remasking, where high-confidence background tokens are unmasked early. The third row (APS-I) adds quantized expectation, which mitigates sampling bias and improves consistency with the measurement, but still relies on prior-based confidence remasking. This results in blurry reconstructions. Finally, the fourth row (APS-II) combines quantized expectation with anchored remasking, preserving optimized anchor tokens while remasking uninformative background tokens. This combination yields the most stable and measurement-consistent generations, as discussed in §B.3. Indeed, we note that best case one could hope for is the direct VQ-VAE reconstruction (top row, second from left) of the Ground Truth image (top row, leftmost). We observe that APS (top row, rightmost) is comparable in quality to the VQ-VAE reconstruction.
B.3.1 Effect of Standard Remasking
Standard remasking in MaskGIT (Chang et al., 2022) and MMaDA (Yang et al., 2025) proceeds as follows. At each iteration, the base denoiser produces a distribution over all tokens ${\mathbf{x}}_{\theta}({\mathbf{z}}_{t(i)})=\{{\mathbf{x}}^{l}_{\theta}({\mathbf{z}}_{t(i)})\}_{l=1}^{L}$ given the current partially unmasked sequence ${\mathbf{z}}_{t}$. Sampled tokens ${\mathbf{x}}=\{{\mathbf{x}}^{l}\}_{l=1}^{L}$ are drawn independently at each position, and previously unmasked tokens are carried over to the next state ${\mathbf{z}}_{s}$. For masked positions, the confidence of each sampled token ${\mathbf{x}}^{l}$ is defined as
@@ B.3.1 Effect of Standard Remasking
B.3.2 Effect of Quantized Expectation
@@
To address this, we propose quantized expectation3.2.1). Instead of sampling, we tilt the unconditional distribution ${\mathbf{x}}_{\theta}({\mathbf{z}}_{t})$ towards the measurement likelihood, obtaining an approximate posterior ${\mathbf{x}}_{\varphi}({\mathbf{z}}_{t})$. We then optimize in the span of codebook embeddings by treating the tilted probabilities as coefficients in a linear combination and passing their expectation through the decoder using a straight-through estimator. The resulting embedding is then quantized back to the nearest valid token. This procedure implicitly maximizes the measurement likelihood, avoids sampling noise, and enables the discovery of tokens with zero prior probability mass but strong measurement consistency–leading to a better posterior sample.
@@
B.3.3 Effect of Anchored Remasking
@@
B.4 Evaluation Under Identical Discrete Diffusion Prior
We compare APS to G2D2 using the official G2D2 codebase (§B.6) on 100 images from the FFHQ validation set. Both methods use an identical compute budget: 100 reverse diffusion steps, each with 30 inner optimization steps. As shown in Table 5, APS consistently improves over G2D2 across all metrics: PSNR improves from $25.46$ to $26.80$, SSIM from $0.717$ to $0.759$, and LPIPS decreases from $0.350$ to $0.310$. Since the generative prior and compute budget are identical, these performance gains arise purely from our algorithmic innovations: quantized expectation (§3.2.1), anchored remasking (§3.2.2), and the use of a perceptual loss. This experiment demonstrates that APS is not only effective but also compatible with both mask-based and uniform discrete diffusion frameworks.
@@ B.4 Evaluation Under Identical Discrete Diffusion Prior
B.5 Computational Complexity

### v2

@@ C.1.2 Reference-based Stylization
For each reverse step, we perform $100$ optimization steps using an Adam optimizer with a learning rate of $0.1$. Gradients from this style loss are backpropagated through the frozen VQ-VAE decoder ${\mathcal{D}}(\cdot)$ and the straight-through estimator $\tilde{{\mathbf{x}}}$ to update all the entries of the conditional probability table $\varphi_{t(i)}=\{\varphi_{t(i)}^{l}\}_{l=1}^{L}$.
In the second stage (§3.2), the posterior estimate ${\mathbf{x}}_{\varphi}({\mathbf{z}}_{t(i)})$ obtained from the first stage is used adaptively unmask anchor tokens in the sequence. Both the processes continue over $15$ total steps following a cosine mask schedule. The fully unmasked sequence satisfies the stylistic constraints without requiring any task-specific retraining of the foundation model.
C.1.3 High-Resolution Inference
@@
C.1.4 Question Answering
In standard discrete diffusion frameworks, the model predicts logits for every position in the sequence, from which tokens are sampled independently. For images, these tokens are in the latent space and hence projected via a decoder into pixel space. In that setting, the combination of the decoder and a measurement error term effectively acts as a proxy for a reward model, allowing our approach APS to provide gradient-like guidance.
In the context of diffusion language modeling, we replace the combination of decoder and measurement error with a scalar reward model (e.g., Qwen3-0.6B) that evaluates the quality of the generated text. The input to this reward model is discrete. The non-differentiable nature of the sampling step breaks the computation graph, preventing the direct backpropagation of the reward signal to the diffusion logits.
Our proposed method, APS, addresses this intractability. APS derives gradient-like guidance through quantized expectation, allowing us to “tilt” the logits of the base model in the direction of higher reward. This enables effective guidance through the discrete bottleneck, analogous to how gradients guide the generation process in discrete image domains.
We integrate our APS sampler into the inference pipeline of Dream-7B-Instruct For all question answering experiments, we fix the total diffusion steps at $T=128$ and the maximum generation length at 128 tokens. We perform $M=3$ gradient optimization steps per diffusion timestep with a learning rate of $\eta=0.1$ to update the intermediate logits. During inference, we generate tokens using a sampling temperature of $0.2$ and nucleus sampling with $p=0.95$. Finally, the remasking temperature is set to $0.0$, ensuring our underlying masking schedule remains consistent with the baseline.
C.2 Algorithm Details
Algorithm 1 Test-Time Anchored Posterior Sampling (APS)
1:Input: Denoising steps $T$, measurement $\mathbf{y}$}, denoiser ${\mathbf{x}}_{\theta}^{\mathrm{logits}}(\cdot)$, operator ${\mathcal{A}}(\cdot)$, decoder ${\mathcal{D}}(\cdot)$.
@@ C.2 Algorithm Details
5:for $i=T$ to 11 do
6:  $\varphi\leftarrow{\mathbf{x}}_{\theta}^{\mathrm{logits}}({\mathbf{z}}_{t(i)})$ $\triangleright$ Quantized Expectation §3.1
7:  for $m=1$ to $M$ do
8:   ${\mathbf{x}}_{\varphi}({\mathbf{z}}_{t(i)})=\mathrm{Softmax}(\varphi)$
9:   $\bar{{\mathbf{x}}}\leftarrow\sum_{k=1}^{K}{\mathbf{c}}_{k}\cdot{\mathbf{x}}_{\varphi}({\mathbf{z}}_{t(i)})[k]$
10:   ${\mathbf{x}}={\mathcal{Q}}_{\mathrm{lfq}}(\bar{{\mathbf{x}}})$
11:   $\tilde{{\mathbf{x}}}\leftarrow\bar{{\mathbf{x}}}+\big[{\mathbf{x}}-\bar{{\mathbf{x}}}\big]_{\mathrm{sg}}$
@@
13:   $\mathcal{L}\leftarrow\mathcal{L}_{\mathrm{reconstruction}}({\mathcal{A}}(\hat{{\mathbf{x}}}),{\mathbf{y}})+\lambda_{\mathrm{p}}\mathcal{L}_{\mathrm{perceptual}}({\mathcal{A}}(\hat{{\mathbf{x}}}),{\mathbf{y}})$
@@
15:  end for
16:  ${\mathbf{x}}_{\varphi}^{*}({\mathbf{z}}_{t(i)})=\mathrm{Softmax}(\varphi)$ $\triangleright$ Anchored Remasking §3.2
17:  $\bar{{\mathbf{x}}}^{*}\leftarrow\sum_{k=1}^{K}{\mathbf{c}}_{k}\cdot{\mathbf{x}}_{\varphi}^{*}({\mathbf{z}}_{t(i)})[k]$
18:  ${\mathbf{x}}={\mathcal{Q}}_{\mathrm{lfq}}(\bar{{\mathbf{x}}}^{*})$
19:  $\kappa^{l}\leftarrow\langle({\mathbf{x}}_{\varphi}^{*}({\mathbf{z}}_{t(i)})^{l},{\mathbf{x}}^{l}\rangle$ for $l=1,\dots,L$
20:  ${\mathcal{P}}_{t(i)}\leftarrow\{l:\kappa^{l}\geq\tau_{t(i)}\}$
21:  ${\mathbf{z}}_{s(i)}\leftarrow\text{Update State}({\mathbf{z}}_{t(i)},{\mathbf{x}},{\mathcal{P}}_{t(i)})$
22:end for
@@
24:return $\hat{{\mathbf{x}}}_{\mathrm{final}}$
Algorithm 1 details our APS procedure. The process begins with a fully masked latent space, ${\mathbf{z}}_{1}$, and iteratively refines the image over $T$ reverse diffusion steps. Each step features an inner optimization phase designed to align the model’s predictions with the given measurement $\mathbf{y}$}. To enable gradient-based optimization through the discrete quantization step which assigns each dimension of an embedding to its nearest value in $\{-1,1\}$ we employ the straight-through estimator (STE) via the stop-gradient operator, $\text{sg}(\cdot)$.
This optimization is guided by a composite loss function. For reconstruction, we primarily use the Mean Absolute Error (MAE, or L1 loss), which we find produces perceptually superior results compared to the Mean Squared Error (MSE). While MSE corresponds to maximizing the Gaussian log-likelihood, MAE is typically more robust. This is supplemented with a VGG perceptual loss to enforce similarity in the feature domain, further improving visual quality. Finally, a small entropy term is included to regularize the posterior distribution. Once the optimization at a given step is complete, the Anchored Remasking strategy uses the model’s confidence defined as the probability assigned to the chosen token for each position to selectively unmask tokens for the next iteration.
@@ C.3 Ablation Study
Figure 6: Ablation study of design choices in APS. We study the posterior sampling problem with the input of the noisy image of a bird (top row, third image from left). Since MMaDA (second row) uses an unconditional prior (no measurement) as the backbone, the final image is inconsistent with the noisy input image (image of person instead of bird). Thus, this shows the effect of standard remasking, where high-confidence background tokens are unmasked early. The third row (APS-I) adds quantized expectation, which mitigates sampling bias and improves consistency with the measurement, but still relies on prior-based confidence remasking. This results in blurry reconstructions. Finally, the fourth row (APS-II) combines quantized expectation with anchored remasking, preserving optimized anchor tokens while remasking uninformative background tokens. This combination yields the most stable and measurement-consistent generations, as discussed in Appendix C.3. Indeed, we note that best case one could hope for is the direct VQ-VAE reconstruction (top row, second from left) of the Ground Truth image (top row, leftmost). We observe that APS (top row, rightmost) is comparable in quality to the VQ-VAE reconstruction.
C.3.1 Effect of Standard Remasking
Standard remasking in MaskGIT (Chang et al., 2022) and MMaDA (Yang et al., 2025b) proceeds as follows. At each iteration, the base denoiser produces a distribution over all tokens ${\mathbf{x}}_{\theta}({\mathbf{z}}_{t(i)})=\{{\mathbf{x}}^{l}_{\theta}({\mathbf{z}}_{t(i)})\}_{l=1}^{L}$ given the current partially unmasked sequence ${\mathbf{z}}_{t}$. Sampled tokens ${\mathbf{x}}=\{{\mathbf{x}}^{l}\}_{l=1}^{L}$ are drawn independently at each position, and previously unmasked tokens are carried over to the next state ${\mathbf{z}}_{s}$. For masked positions, the confidence of each sampled token ${\mathbf{x}}^{l}$ is defined as
@@ C.3.1 Effect of Standard Remasking
C.3.2 Effect of Quantized Expectation
@@
To address this, we propose quantized expectation3.1). Instead of sampling, we tilt the unconditional distribution ${\mathbf{x}}_{\theta}({\mathbf{z}}_{t})$ towards the measurement likelihood, obtaining an approximate posterior ${\mathbf{x}}_{\varphi}({\mathbf{z}}_{t})$. We then optimize in the span of codebook embeddings by treating the tilted probabilities as coefficients in a linear combination and passing their expectation through the decoder using a straight-through estimator. The resulting embedding is then quantized back to the nearest valid token. This procedure implicitly maximizes the measurement likelihood, avoids sampling noise, and enables the discovery of tokens with zero prior probability mass but strong measurement consistency–leading to a better posterior sample.
@@
C.3.3 Effect of Anchored Remasking
@@
[…]
# Optimization Steps ($M$) PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
0 9.49 0.1855 0.6747
10 21.47 0.5541 0.4298
50 24.80 0.7130 0.2730
100 25.71 0.7422 0.2468
200 25.72 0.7389 0.2474
C.3.4 Effect of Optimization Steps
Table 6 presents an ablation study on the number of inner optimization steps used during test-time anchoring in Algorithm 1. We observe that performance improves rapidly up to 50 steps and saturates beyond 100 steps, confirming that our method converges efficiently without requiring additional iterations. Importantly, all optimization steps backpropagate only through the lightweight VQ-VAE decoder rather than the large 8B-parameter MMaDA backbone, keeping computational overhead minimal.
C.4 Evaluation Under Identical Discrete Diffusion Prior
We compare APS to G2D2 using the official G2D2 codebase (§C.7) on 100 images from the FFHQ validation set. Both methods use an identical compute budget: 100 reverse diffusion steps, each with 30 inner optimization steps. As shown in Table 7, APS consistently improves over G2D2 across all metrics: PSNR improves from $25.46$ to $26.80$, SSIM from $0.717$ to $0.759$, and LPIPS decreases from $0.350$ to $0.310$. Since the generative prior and compute budget are identical, these performance gains arise purely from our algorithmic innovations: quantized expectation (§3.1), anchored remasking (§3.2), and the use of a perceptual loss. This experiment demonstrates that APS is not only effective but also compatible with both mask-based and uniform discrete diffusion frameworks.
@@ C.4 Evaluation Under Identical Discrete Diffusion Prior
C.5 Computational Complexity
C.5.1 Tokenwise Expectation to Differentiable Surrogate
In this section, we provide a detailed derivation of our optimization objective (3.1) from Equation (3.1), and analyze the associated computational complexity.
Define the measurement term in Equation (3.1) as
${\mathcal{L}}_{t}^{\text{meas}}~:=~\mathbb{E}_{q(Z_{s}|{\mathbf{z}}_{t},{\mathbf{x}})}\Big[\sum_{l=1}^{L}-\log q\big({\mathbf{y}}\,\big|\,{\mathbf{x}}_{\varphi}({\mathbf{z}}_{t};Z_{s}^{l})\big)\Big]=\sum_{l=1}^{L}\mathbb{E}_{q(Z_{s}^{l}|{\mathbf{z}}_{t}^{l},{\mathbf{x}})}\!\Big[-\log q\big({\mathbf{y}}\,\big|\,{\mathbf{x}}_{\varphi}({\mathbf{z}}_{t};Z_{s}^{l})\big)\Big],$ (21)
where the last equality is due to $q(Z_{s}|{\mathbf{z}}_{t},{\mathbf{x}})$ factorizing across sites given ${\mathbf{z}}_{t}$ (the forward process is conditionally independent tokenwise). Each inner expectation depends on the local categorical $q(Z_{s}^{l}|{\mathbf{z}}_{t}^{l},{\mathbf{x}})$, which (under the mask process) has support on only two outcomes (e.g., $\mathbf{x}$} or $\mathbf{m}$}). Evaluating ${\mathcal{L}}_{t}^{\text{meas}}$ exactly would require propagating each choice through $\mathcal{D}$} and $\mathcal{A}$}, which is expensive in practice.
Complexity of the exact marginalization. We emphasize that evaluating the measurement term exactly is exponential in the sequence length $L$. Concretely, the outer marginalization over the tokenwise mask choices $Z_{s}$ involves at most $2^{L}$ joint assignments (each site is either $\mathbf{m}$} or set to $\mathbf{x}$} under the mask process). For each fixed assignment $Z_{s}$, computing the inner expectation $\mathbb{E}_{X\sim{\mathbf{x}}_{\varphi}({\mathbf{z}}_{t}),\,X^{l}\leftarrow Z_{s}^{l}}[-\log q({\mathbf{y}}|X)]$ requires aggregating the likelihood over all possible completions drawn accordingly. If the decoder has a discrete codebook of size $K$ per position, the number of such completions is upper-bounded by $K^{L}$ (or more tightly $K^{L-1}$ since $l$-the position is fixed with $Z_{s}^{l}$), so a worst-case cost for the inner expectation is $O(K^{L})$. Combining these two sources of combinatorial explosion gives a conservative worst-case cost of $O(2^{L}\cdot K^{L})=O\big((2K)^{L}\big)$, i.e., exponential in $L$. Thus, while one can write down the exact marginalization formally, it is computationally intractable for realistic sequence lengths and vocabulary sizes.
Complexity of the proposed approximation. We adopt a plug-in approximation that replaces the tokenwise expectation by a single differentiable surrogate. Let $h({\mathbf{x}}):=-\log q({\mathbf{y}}|{\mathbf{x}})$ and $h({\mathbf{x}}_{\varphi}({\mathbf{z}}_{t};{\mathbf{z}}_{s}^{l}))$ be defined as in §3, which we recall here for clarity: $h({\mathbf{x}}_{\varphi}({\mathbf{z}}_{t};{\mathbf{z}}_{s}^{l}))=-\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}({\mathbf{z}}_{t};{\mathbf{z}}_{s}^{l}))=-\mathbb{E}_{X\sim{\mathbf{x}}_{\varphi}({\mathbf{z}}_{t}),X^{l}\leftarrow{\mathbf{z}}_{s}^{l}}[\,\log q({\mathbf{y}}|X)\,]$. Then, a first-order Taylor’s expansion of each term inside the summation of (21) around ${\mathbf{u}}\coloneqq\mathbb{E}_{Z_{s}^{l}\sim q(\cdot|{\mathbf{z}}_{t}^{l},{\mathbf{x}})}\mathbb{E}_{X\sim{\mathbf{x}}_{\varphi}({\mathbf{z}}_{t}),X^{l}\leftarrow Z_{s}^{l}}\Big[X\Big]$ yields:
$\mathbb{E}_{Z_{s}^{l}\sim q(\cdot|{\mathbf{z}}_{t}^{l},{\mathbf{x}})}\Big[h\big({\mathbf{x}}_{\varphi}({\mathbf{z}}_{t};Z_{s}^{l})\big)\Big]$ $=\mathbb{E}_{Z_{s}^{l}\sim q(\cdot|{\mathbf{z}}_{t}^{l},{\mathbf{x}})}\Big[-\log q\big({\mathbf{y}}|{\mathbf{x}}_{\varphi}({\mathbf{z}}_{t};Z_{s}^{l})\big)\Big]=\mathbb{E}_{Z_{s}^{l}\sim q(\cdot|{\mathbf{z}}_{t}^{l},{\mathbf{x}})}\mathbb{E}_{X\sim{\mathbf{x}}_{\varphi}({\mathbf{z}}_{t}),X^{l}\leftarrow Z_{s}^{l}}\Big[-\log q\big({\mathbf{y}}|X\big)\Big]$
$\approx\mathbb{E}_{Z_{s}^{l}\sim q(\cdot|{\mathbf{z}}_{t}^{l},{\mathbf{x}})}\mathbb{E}_{X\sim{\mathbf{x}}_{\varphi}({\mathbf{z}}_{t}),X^{l}\leftarrow Z_{s}^{l}}\Big[-\log q\big({\mathbf{y}}|{\mathbf{u}}\big)-\langle\nabla\log q({\mathbf{y}}|{\mathbf{u}}),X-{\mathbf{u}}\rangle\Big]$
(Ignore higher order terms in the Taylor’s series expansion.)
$=-\log q\big({\mathbf{y}}|{\mathbf{u}}\big)-\langle\nabla\log q({\mathbf{y}}|{\mathbf{u}}),~\mathbb{E}_{Z_{s}^{l}\sim q(\cdot|{\mathbf{z}}_{t}^{l},{\mathbf{x}})}\mathbb{E}_{X\sim{\mathbf{x}}_{\varphi}({\mathbf{z}}_{t}),X^{l}\leftarrow Z_{s}^{l}}\Big[X\Big]-{\mathbf{u}}\rangle$
[…]
$=h\Big(\mathbb{E}_{Z_{s}^{l}\sim q(\cdot|{\mathbf{z}}_{t}^{l},{\mathbf{x}})}\mathbb{E}_{X\sim{\mathbf{x}}_{\varphi}({\mathbf{z}}_{t}),X^{l}\leftarrow Z_{s}^{l}}\Big[X\Big]\Big).$
With this approximation, we avoid the computation of $\log q\big({\mathbf{y}}|X\big)$ for every sample $X$. This would otherwise require a function call through ${\mathcal{A}}\!\circ\!{\mathcal{D}}$ per sample $X$, which is expensive. In contrast, the double expectation inside $h(\cdot)$ scales linearly in sequence length ($L$) and vocabulary size ($K$) because for each position of the sequence $X=(X^{1},\ldots,X^{L})$, the expectation can be independently computed using ${\mathbf{x}}^{l}_{\varphi}({\mathbf{z}}_{t})$ that costs ${\mathcal{O}}(K)$, totaling up to ${\mathcal{O}}(KL)$. With the outer sum over $L$ tokens in (21), the total complexity of ${\mathcal{L}}_{t}^{\text{meas}}$ becomes ${\mathcal{O}}(KL^{2})$.
Next, we discuss how to reduce the complexity from ${\mathcal{O}}(KL^{2})$ to ${\mathcal{O}}(KL)$. Note that ${\mathbf{x}}_{\varphi}({\mathbf{z}}_{t};Z_{s}^{l})$ differs from ${\mathbf{x}}_{\varphi}({\mathbf{z}}_{t})$ only at position $l$. Assuming local smoothness of ${\mathcal{A}}\!\circ\!{\mathcal{D}}$, we approximate $q(\cdot|{\mathbf{z}}_{t}^{l},{\mathbf{x}})$ by ${\mathbf{x}}^{l}_{\varphi}({\mathbf{z}}_{t})$ resulting in the quantized expectation: for each site, form the expected codebook embedding $\bar{{\mathbf{x}}}^{l}=\sum_{k}{\mathbf{c}}_{k}\,{\mathbf{x}}_{\varphi}^{l}({\mathbf{z}}_{t})[k]$, quantize via LFQ as ${\mathbf{x}}^{l}={\mathcal{Q}}_{\text{lfq}}(\bar{{\mathbf{x}}}^{l})$, and apply the straight-through estimator to obtain a single surrogate sequence $\tilde{{\mathbf{x}}}_{\varphi}({\mathbf{z}}_{t})=\bar{{\mathbf{x}}}+[{\mathbf{x}}-\bar{{\mathbf{x}}}]_{\text{sg}}$. Substituting the same surrogate for all $l$ gives the following [^6] In practice, the factor $L$ can be absorbed into the likelihood weight or learning rate. If only positions with ${\mathbf{z}}_{t}^{l}={\mathbf{m}}$ contribute, one may replace $L$ by $|\{l:{\mathbf{z}}_{t}^{l}={\mathbf{m}}\}|$. We explicitly show the absorbed constant here for clarity.:
${\mathcal{L}}_{t}^{\text{meas}}~\approx~\sum_{l=1}^{L}h\!\big(\tilde{{\mathbf{x}}}_{\varphi}({\mathbf{z}}_{t})\big)~=~L\cdot\big(-\log q({\mathbf{y}}|\tilde{{\mathbf{x}}}_{\varphi}({\mathbf{z}}_{t}))\big).$